OpenStudy (anonymous):

What are the solutions of 2x2 + 8x = -26 A) -2 ± 3 i B) -2 ± i C) 2 ± 6 i D) -3 ± 2 i

OpenStudy (anonymous):

Write the equation as \[ 2 x^2 +8 x +26 =0 \] and compute \[ b^2-4ac \]

OpenStudy (anonymous):

You can simplify it by dividing by 2 to get \[ x^2 + 4x +13=0 \]

OpenStudy (anonymous):

so 8^2-4(2)(-46)?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

Use the simplified version \[ x^2 + 4x +13=0\\ a=1; b =4; c=13 \]

OpenStudy (anonymous):

\[ b^2 - 4 a c = 4^2 - 4(1)13=16- 52=-36 \]

OpenStudy (anonymous):

\[ \frac {- 4 \pm i \sqrt {36}}{2}= -2 \pm 3i \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

So it is A